The product of 1 and double a number is 7. Find the number.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

number is x
2x=7
x=3.5

Answer by ikleyn(52908)

.

    1*(2x) = 7


    2x = 7


    x = 7/2 = 3.5.    ANSWER

Solved and explained.